A solution contains 7.1 parts per million oxygen. There are 0.027 grams of oxygen dissolved in the solution. How much of the solvent water in grams are in the solution
1
Expert's answer
2020-04-13T02:26:06-0400

The number or moles of oxygen O2 is 0.027 g / 32 g/mole = 8.44*10-4 mole

The number of moles of water is that of oxygen divided by concentration:

8.44*10-4 / 7.1*10-6 = 118.87 mole

the mass of water H2O is 118.87 mole * 18 g/mole = 3400 g